# Science Out of Its Ivory Tower: Improving Accessibility with Reinforcement Learning

### v1

@@ 6.3 Token Distribution Shift Analysis for Reinforcement Learning Models
We also analyzed the distribution of token types with respect to their positions in the generated sequences. In Figure 3, the top two subplots show the average distributions from the four RLAM models for clarity. The left column focuses on marginal tokens, while the right column visualizes shifted tokens. Each subplot illustrates the proportion of each token category relative to the total token count at each position. Our analysis reveals that the RLAM models exhibit more marginal modifications at the initial positions, whereas shifted tokens tend to accumulate toward the end of the sequences. This pattern suggests that, compared to the supervised fine-tuning model, which tends to overemphasize implications (Wang and Clark, 2024), the RLAM models may truncate redundant content by employing more shifted tokens toward the end. Marginal shifts are used at the beginning to introduce main subjects or background information with simpler words.
In contrast, the RLARI model does not display such a systematic pattern; different types of tokens appear to be more evenly distributed across positions but lack the level of consistency observed in the RLAM models. We hypothesize that this difference may be caused by disproportionately large incentives for shorter sentences, which could hinder effective exploration of alternative, simpler word substitutions and ultimately contribute to the RLARI model’s performance variability after the initial phases of optimization. Additionally, approximately 90 percent of the tokens generated by the reinforcement learning models remain unchanged, which roughly corroborates previous findings in RLHF studies, where around 8 percent are changed (Lin et al., 2023). The difference may arise from the nature of the task; we focus on rephrasing in an accessible manner, which often does not require altering many tokens, unlike Lin et al. (2023), who align language models to human preferences, where the expected answer can be quite different from what would otherwise be derived from the base model.
@@ 7 Conclusion
We considered the limitations of using average sentence length as a reward, as it can be a rather coarse measurement of syntactic complexity and perhaps also of cognitive comprehensibility. We observed that it is not uncommon for restrictive attributive clauses and other syntactic structures to appear in the generations from RLAM-trained models that could theoretically be split into separate sentences. However, increasing the coefficient for the sentence length reward exacerbated optimization instability, as was observed in the RLARI models. In future work, more direct yet easily obtainable measures, such as dependency length (Futrell et al., 2015), could be a promising direction to explore for generating shorter sentences.
Another potential direction is to generate lay summaries from the full text, not just the abstract. With current large language models capable of handling contexts with several thousand tokens, potentially enhanced by rotary positional encoding (Su et al., 2024), this approach is feasible. This method also has the added benefit of reducing hallucinations by providing more comprehensive information from the entire manuscript. Also, reinforcement learning from accessibility measures is not limited to scientific literature; it can also be an effective solution for other simplification scenarios where a gold-standard corpus is unavailable. We hope this work contributes to bridging the gap between scholarship and a broader audience, advancing the understanding and development of better simplification systems, and ultimately fostering a more informed and engaged society.
@@ Appendix A Reinforcement Learning Using PPO
We used the PPO (Schulman et al., 2017) clipped surrogate objective with importance sampling to more efficiently use offline samples to update the online policy. Importance sampling corrects for the discrepancy between the behavior policy that generated the samples and the current policy by weighting the samples using the ratio of their probabilities under both policies. The PPO algorithm introduces a clipping mechanism to balance exploration and exploitation while preventing large, potentially harmful updates to the policy (see Equation 4).
@@ Appendix B Adaptive KL Controller
Following Ziegler et al. (2019, Sec. 2.2), we dynamically adjust $\beta_{\text{KL}}$ to target a specific KL divergence value, $\text{KL}_{\text{target}}$, using a log-space proportional controller.

### v2

@@ 6.3 Token Distribution Shift Analysis for Reinforcement Learning Models
We also analyzed the distribution of token types with respect to their positions in the generated sequences. In Figure 4, the top two subplots show the average distributions from the four RLAM models for clarity. The left column focuses on marginal tokens, while the right column visualizes shifted tokens. Each subplot illustrates the proportion of each token category relative to the total token count at each position. Our analysis reveals that the RLAM models exhibit more marginal modifications at the initial positions, whereas shifted tokens tend to accumulate toward the end of the sequences. This pattern suggests that, compared to the supervised fine-tuning model, which tends to overemphasize implications (Wang and Clark,, 2025), the RLAM models may truncate redundant content by employing more shifted tokens toward the end. Marginal shifts are used at the beginning to introduce main subjects or background information with simpler words.
In contrast, the RLARI model does not display such a systematic pattern; different types of tokens appear to be more evenly distributed across positions but lack the level of consistency observed in the RLAM models. We hypothesize that this difference may be caused by disproportionately large incentives for shorter sentences, which could hinder effective exploration of alternative, simpler word substitutions and ultimately contribute to the RLARI model’s performance variability after the initial phases of optimization. Additionally, approximately 90 percent of the tokens generated by the reinforcement learning models remain unchanged, which roughly corroborates previous findings in RLHF studies, where around 8 percent are changed (Lin et al.,, 2023). The difference may arise from the nature of the task; we focus on rephrasing in an accessible manner, which often does not require altering many tokens, unlike Lin et al., (2023), who align language models to human preferences, where the expected answer can be quite different from what would otherwise be derived from the base model.
@@ 7 Conclusion
We considered the limitations of using average sentence length as a reward, as it can be a rather coarse measurement of syntactic complexity and perhaps also of cognitive comprehensibility. We observed that it is not uncommon for restrictive attributive clauses and other syntactic structures to appear in the generations from RLAM-trained models that could theoretically be split into separate sentences. However, increasing the coefficient for the sentence length reward exacerbated optimization instability, as was observed in the RLARI models. In future work, more direct yet easily obtainable measures, such as dependency length (Futrell et al.,, 2015), could be a promising direction to explore for generating shorter sentences.
Another potential direction is to generate lay summaries from the full text, not just the abstract. With current large language models capable of handling contexts with several thousand tokens, potentially enhanced by rotary positional encoding (Su et al.,, 2024), this approach is feasible. This method also has the added benefit of reducing hallucinations by providing more comprehensive information from the entire manuscript. Also, reinforcement learning from accessibility measures is not limited to scientific literature; it can also be an effective solution for other simplification scenarios where a gold-standard corpus is unavailable. We hope this work contributes to bridging the gap between scholarship and a broader audience, advancing the understanding and development of better simplification systems, and ultimately fostering a more informed and engaged society.
@@ Appendix A Reinforcement Learning Using PPO
We used the PPO (Schulman et al.,, 2017) clipped surrogate objective with importance sampling to more efficiently use offline samples to update the online policy. Importance sampling corrects for the discrepancy between the behavior policy that generated the samples and the current policy by weighting the samples using the ratio of their probabilities under both policies. The PPO algorithm introduces a clipping mechanism to balance exploration and exploitation while preventing large, potentially harmful updates to the policy (see Equation 4).
@@ Appendix B Adaptive KL Controller
Following Ziegler et al., (2019, Sec. 2.2), we dynamically adjust $\beta_{\text{KL}}$ to target a specific KL divergence value, $\text{KL}_{\text{target}}$, using a log-space proportional controller.
@@
Appendix C Manual Evaluation
The evaluation involved three annotators: Jason Clark (academic librarian), Zuoyu Tian (computational linguist), and Haining Wang (information scientist). The following rubrics were iteratively developed and refined over two months of model monitoring and interim generation reviews.
  • Language Quality: Evaluates the fluency and grammatical accuracy of the text.
    • Good: Natural and free from grammatical errors; expressions flow properly.
    • Acceptable: Natural and error-free, although it may contain formulaic endings (e.g., “All rights reserved.”) that have minimal impact on readability.
    • Poor: Contains grammatical errors.
  • Completeness: Assesses the coverage of the abstract’s main components, including context, purpose, main findings, and conclusions.
    • Good: Fully covers all four aspects without truncation.
    • Acceptable: Covers at least three aspects, though the generation may be truncated.
    • Poor: Covers fewer than three of the main aspects.
  • Faithfulness: Assesses the scientific fidelity of the simplified text to the original abstract.
    • Good: Accurately reflects the main claims of the original abstract, with no factual errors.
    • Acceptable: Maintains semantic alignment, though the claims remain unverifiable after 15 minutes of manual searching without AI assistance. No factual errors are present.
    • Poor: Misrepresents findings, contains false claims, or includes factual errors.
Each annotator independently reviewed the selected samples using the finalized rubrics before beginning formal annotation. In the formal rating phase, discrepancies were resolved through discussion, referencing the original manuscripts and external sources as needed. While our evaluations of language quality and completeness were generally consistent, assessing faithfulness proved challenging given the broad disciplinary diversity of the source abstracts. For example, the original sentence states: “At high concentrations, Zn2+, in addition, binds to a second site and inhibits the outward movement of the voltage sensor of Hv1.” A simplified version reads: “At high levels, it also prevents the channel’s parts from moving.” While the simplified version is notably more accessible and preserves the core causal relationship, it remains unclear—without specialized domain expertise—whether the omission of “outward” specifies a subtle factual distortion (e.g., whether inward movement is also implicated). In these instances, we prioritized semantic alignment for a general audience. In retrospect, we acknowledge that incorporating inter-rater agreement metrics, such as Cohen’s Kappa, for language quality and completeness would have enhanced our methodological rigor.
We have made all generations, including intermediate outputs from the RL training process, publicly available in our project repository. We encourage readers and domain experts to evaluate the faithfulness of our outputs in light of their disciplinary expertise.